OpenStudy (anonymous):

6. Find the length of the side of a square with an area of 144. (1 point) • 48 • 12 • 206 • 288

OpenStudy (anonymous):

144/2

OpenStudy (anonymous):

12

OpenStudy (anonymous):

72

OpenStudy (ribhu):

its sqrt of 144 which is 12

OpenStudy (ribhu):

area = side *side

OpenStudy (anonymous):

72 *2 does not =12

OpenStudy (k_lynn):

If you know the area of a square, you cal always find the side by finding the square root. So what's \[\sqrt{144}\]Or what number times itself gets you the number 144?

OpenStudy (anonymous):

12*12= 144! that easy

OpenStudy (k_lynn):

https://learnzillion.com/lessons/187-find-the-side-lengths-of-a-square-given-its-area That link has a video on how to do those kinds of problems :)

OpenStudy (anonymous):

Sorry got mixed up Thanks!

OpenStudy (anonymous):

its 12 definitly

OpenStudy (anonymous):

your welcome